karissafrazier:

Evaluate: { 25/36} 1/2 A. 5/6 B. 5/36 C. 25/36 D. 25/6

jhonyy9:

\[\frac{ 25 }{ 36 } * \frac{ 1 }{ 2 } = ? \] this is your question ?

karissafrazier:

no- the 1/2 is an exponent

karissafrazier:

Rival:

Do you got any notes?

karissafrazier:

no not really

Florisalreadytaken:

\[ \left(\frac{25}{36}\right)^{\frac{1}{2}} \Rightarrow \sqrt{\frac{25}{36} } \Rightarrow \frac{\sqrt{25}}{\sqrt{36}} \] what is \( \sqrt{25} \) and \( \sqrt{36} \) ??

Rival:

25−√ =5 36-√ = 6

karissafrazier:

can i somehow get a reply lock on this?